OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (amistre64):

to write the equation of a line youd need to know its slope; or you can go the long route and just plug in the slope formula where the slope is spose to go :) \[y = \frac{y-y_0}{x-x_0}x -\frac{(y-y_0)(x_0)}{x-x_0}+y_0\]

OpenStudy (amistre64):

which form of the equation of a line are you looking to write?

OpenStudy (anonymous):

Slope Intercept form or Point Slope Form.

OpenStudy (amistre64):

point slope form would be the easier construction then; (y-y0) = m(x-x0); where P(x,y) and P(x0,y0) are the 2 points given

OpenStudy (amistre64):

and m= (y-y0)/(x-x0)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

O but wait, Why did you choose the point you did?

OpenStudy (amistre64):

It doesnt matter which point you choose to go where. they both amount to the same thing in the end

OpenStudy (amistre64):

(–1, –4) and (2, 5) y-2=9/2 (x-5) is the same as y+1=9/2(x+4) ... if I got the slope right

OpenStudy (kirbykirby):

Find the slope of your line: (y2 - y1)/(x2 - x1) = (-4 - 5)/(-1 - 2) = -9/-3 = 3 Now you have this: y = 3x + b. In order to find the intercept, just plug in either given point: Using (2, 5): we have 5 = 3(2) + b 5 = 6 + b -1 = b It doesn't matter which point you use. Both points are on the SAME line, so both equations will satisfy he equation (You can try using (-1, -4) if you want to be sure). Why choose (2, 5)? Well you could say you don't like dealing with negative numbers? The second part of the question is rather stupid.

OpenStudy (amistre64):

I think my original post confused the slope formula in the process :)

OpenStudy (anonymous):

No Problem. Thank you both.